OpenStudy (anonymous):

An equation is shown below. 2x + y = 3 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points) HELP PLEASE

OpenStudy (anonymous):

@Lena772 @SolomonZelman @mathmale @undeadknight26 @seventhgradealgebra @oleg3321

OpenStudy (anonymous):

@gswag98

OpenStudy (anonymous):

@thecatmike

OpenStudy (anonymous):

@jamieyang08

OpenStudy (anonymous):

FOR PART A: 1 solution would be like subtracting "2x" from both sides

OpenStudy (anonymous):

so then its y =3-2x?

OpenStudy (anonymous):

no like this y = -2x + 3

OpenStudy (anonymous):

O ok so then I need more solutions right?

OpenStudy (anonymous):

another solution would be like y being part or unknown of in another equation

OpenStudy (anonymous):

like what ?

OpenStudy (anonymous):

like maybe 2x + 2y = 40 or something

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and then you plug in what "y" equals then solve

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

soo @gswag98 whats b now I need 3 different parts?

OpenStudy (anonymous):

well you can use the two that i gave you for Part A maybe

OpenStudy (anonymous):

do i plug in for x and y

OpenStudy (anonymous):

whatcha mean??

OpenStudy (anonymous):

can i plug something in for x and y @gswag98

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) the answer could be: "plot the line defined by this equation. All points on the line are (x,y) pairs which are solutions to the equation"

OpenStudy (anonymous):

So its not asking me for a equation.?

OpenStudy (phi):

to plot the line, you can change from standard form to slope-intercept form (but you don't have to) However, to find 3 points on the line, it's helpful to rewrite the equation as y = -2x+3 (btw you could write y= 3-2x, but it's convention to write the "x" part first)

OpenStudy (phi):

to find 3 points, pick "nice numbers" for x, replace the x with that number, and find y

OpenStudy (anonymous):

what do you mean nice numbers

OpenStudy (phi):

easy to do the arithmetic. example: x=0 is easy to do. 3.14159 is not so easy

OpenStudy (anonymous):

Oh yea haha

OpenStudy (anonymous):

so i will plug in 2

OpenStudy (anonymous):

an it =-1

OpenStudy (phi):

that means when x=2, y = -1 (2,-1) is a point on the line, and a "solution" if you use x=2 and y=-1 in the original equation, you will see it "works"

OpenStudy (anonymous):

O yesssss hold on let me write this down

OpenStudy (anonymous):

Ok so I have to make 2 more of those right?

OpenStudy (phi):

yes. what did you come up with?

OpenStudy (anonymous):

I came up with 3 and 5 i will do

OpenStudy (anonymous):

i meant -5

OpenStudy (anonymous):

I dont understand the last part @phi

OpenStudy (phi):

First, what points did you get for part B?

OpenStudy (phi):

For part C, they want you to write another equation to form a system of equations a system of equations means the solution must work for *both* equations at the same time. Because your equations are lines, it means the solution must be on *both lines*. That only happens where the lines cross. At the intersection point, that point is on both lines.

OpenStudy (anonymous):

so they have to intersect like?

OpenStudy (phi):

system of equations that is inconsistent INCONSISTENT means no solution no solution means the lines never cross. lines that never cross are PARALLEL so you need to find the equation of a line parallel to the original. Can you do that ?

OpenStudy (anonymous):

yes I understand

OpenStudy (anonymous):

-2x-3

OpenStudy (anonymous):

Thank you for your help!!!!!

OpenStudy (phi):

if you put y= in front of it

OpenStudy (anonymous):

yess